# QFNN-FFD: Quantum Federated Neural Network for Financial Fraud Detection

### v3

@@ I Introduction
In the rapidly evolving financial technology landscape, privacy is a fundamental pillar, crucial for upholding the trust and integrity of financial transactions and services [1]. As digital transactions become more prevalent, the volume of sensitive data handled by financial institutions grows exponentially, making robust privacy measures indispensable [2]. The emergence of Quantum Machine Learning (QML) marks a transformative era [3, 4, 5, 6, 7, 8, 9], promising computational capabilities by exploiting quantum physics [10], while simultaneously raising pivotal concerns about privacy and data security. This paper introduces the Quantum Federated Neural Network for Financial Fraud Detection (QFNN-FFD), a framework that integrates the quantum-enhanced processing power of Quantum Computing (QC) with the privacy-preserving attributes of Federated Learning (FL). The synergy of QML with FL jointly improves the efficiency and accuracy of detecting fraudulent activities, while safeguarding sensitive financial data against the ever-looming threats of breaches and unauthorized access.
@@ II Background and Related Works
Advancing into financial fraud, significant research has been conducted to apply QC and QML in detecting financial fraud. In [17], they developed quantum protocols for anomaly detection, applying them to credit card fraud. Furthermore, in [31], they explored using a Quantum Support Vector Machine (QSVM) for real-world financial data, presenting one of the first end-to-end applications of QML in the financial sector. As the application of QML in fraud detection advances, several innovative approaches have emerged. For instance, in [32], they explored using QML models, including the Variational Quantum Classifier (VQC) and different QNNs. These models showed promising results in classifying fraud and non-fraud transactions, demonstrating QML’s potential in financial applications. In [33], the study addressed the latency in traditional fraud detection systems by implementing a QML approach using a Support Vector Machine (SVM) enhanced with quantum annealing solvers. In [34], they discussed a hybrid model that combines QNNs with classical neural networks to enhance fraud detection capabilities.
@@ IV-A Experimental Setup
In our study, we utilize the IEEE-CIS Fraud Detection dataset [44]. It is divided into two primary files: identity and transaction, linked by TransactionID. It encompasses both numerical and categorical features essential for identifying fraudulent activities. The preprocessing steps begin with optimizing the dataset’s memory usage by refining data types, significantly reducing its memory footprint. This is followed by a detailed analysis of missing values, which helps identify and quantify missing data to inform our approach to managing these instances. Subsequently, features are categorized and processed: categorical variables undergo one-hot encoding, while numerical variables are standardized. To counteract the class imbalance between fraud and non-fraud instances, an up-sampling technique is employed to ensure equitable representation of both classes.
Our QFNN-FFD is implemented using PennyLane for model architecture and Qiskit for simulating quantum noise [45], enabling a realistic QC environment. The framework, structured around four qubits, employs the Adam optimizer ($\eta$=0.1) across dual training phases—local and global—with up to 100 iterations for each across 15 clients. This setup is characterized by 32 initially random parameters, which are optimized through evaluations on a training set comprising 115,386 instances (80% of the total dataset of 144,233 instances) and a validation set comprising 28,847 instances, which is 20% of the total dataset. We focus on binary classification accuracy and MSE as key metrics. Operational deployment occurs within an environment characterized by a configuration consisting of 4 virtual CPUs (vCPUs), 25 gigabytes (GB) of RAM, and a single NVIDIA Tesla V100 virtual GPU (vGPU). This setup offers a balance of processing power, memory capacity, and advanced GPU acceleration—crucial factors for efficiently handling the intensive computations required by our QFNN-FFD framework.
@@ IV-B Accuracy and Loss Analysis
IV-C Local Client Validation Accuracy Analysis
IV-D Quantum Noise Analysis
@@ IV-E Comparison with Existing Works
Reference Precision Recall F1-Score Accuracy Privacy [15] 84 84.44 75.68 83.92 $\times$ [16] 96.1 79.5 86 94.5 $\times$ [17] 90 $\times$ Our QFNN-FFD 95 96 95 95
@@
Compared to the results in Table I, our QFNN-FFD outperforms other QML models applied to similar datasets, achieving superior performance metrics. These metrics include precision, recall, F1-score, and accuracy, where QFNN-FFD demonstrates comprehensive superiority across all fronts. This performance is a testament to the model’s efficacy and highlights its ability to effectively integrate complex quantum computations within an FL framework. Unlike the existing models [15, 16, 17], which focus solely on performance, QFNN-FFD additionally integrates a privacy-preserving FL approach. This ensures high detection accuracy of 95% and enhanced data privacy, establishing QFNN-FFD as a leading solution for secure and efficient fraud detection in fintech.
IV-F Discussion

### v5

@@ Abstract
Index Terms:
Quantum Neural Network, Quantum Federated Learning, Quantum Machine Learning, Fraud Detection, Finance, Classification
@@ I Introduction
In the rapidly evolving financial technology landscape, privacy is a fundamental pillar, crucial for upholding the trust and integrity of financial transactions and services [1]. As digital transactions become more prevalent, the volume of sensitive data handled by financial institutions grows exponentially, making robust privacy measures indispensable [2]. The emergence of Quantum Machine Learning (QML) marks a transformative era [3, 4, 5, 6, 7, 8, 9, 10, 11], promising computational capabilities by exploiting quantum physics [12], while simultaneously raising pivotal concerns about privacy and data security. This paper introduces the Quantum Federated Neural Network for Financial Fraud Detection (QFNN-FFD), a framework that integrates the quantum-enhanced processing power of Quantum Computing (QC) with the privacy-preserving attributes of Federated Learning (FL). The synergy of QML with FL jointly improves the efficiency and accuracy of detecting fraudulent activities, while safeguarding sensitive financial data against the ever-looming threats of breaches and unauthorized access.
@@ II Background and Related Works
Advancing into financial fraud, significant research has been conducted to apply QC and QML in detecting financial fraud. In [19], they developed quantum protocols for anomaly detection, applying them to credit card fraud. Furthermore, in [33], they explored using a Quantum Support Vector Machine (QSVM) for real-world financial data, presenting one of the first end-to-end applications of QML in the financial sector. As the application of QML in fraud detection advances, several innovative approaches have emerged. For instance, in [34, 35], they explored using QML models, including the Variational Quantum Classifier (VQC) and different QNNs. These models showed promising results in classifying fraud and non-fraud transactions, demonstrating QML’s potential in financial applications. In [36], the study addressed the latency in traditional fraud detection systems by implementing a QML approach using a Support Vector Machine (SVM) enhanced with quantum annealing solvers. In [37], they discussed a hybrid model that combines QNNs with classical neural networks to enhance fraud detection capabilities.
@@ IV-A Experimental Setup
In our study, we utilize the IEEE-CIS Fraud Detection dataset [47]. It is divided into two primary files: identity and transaction, linked by TransactionID. It encompasses both numerical and categorical features essential for identifying fraudulent activities. The preprocessing steps begin with optimizing the dataset’s memory usage by refining data types, significantly reducing its memory footprint. This is followed by a detailed analysis of missing values, which helps identify and quantify missing data to inform our approach to managing these instances. Subsequently, features are categorized and processed: categorical variables undergo one-hot encoding, while numerical variables are standardized. To counteract the class imbalance between fraud and non-fraud instances, an up-sampling technique is employed to ensure equitable representation of both classes [48].
Our QFNN-FFD is implemented using PennyLane for model architecture and Qiskit for simulating quantum noise [49, 50], enabling a realistic QC environment. We conduct extensive hyperparameter tuning, exploring various configurations before selecting the optimal settings. The framework, structured around four qubits, employs the Adam optimizer ($\eta$=0.1) across dual training phases—local and global—with up to 100 iterations for each across 15 clients. This setup is characterized by 32 initially random parameters, which are optimized through evaluations on a training set comprising 115,386 instances (80% of the total dataset of 144,233 instances) and a validation set comprising 28,847 instances, which is 20% of the total dataset. We focus on binary classification accuracy and MSE as key metrics. Operational deployment occurs within an environment characterized by a configuration consisting of 4 virtual CPUs (vCPUs), 25 gigabytes (GB) of RAM, and a single NVIDIA Tesla V100 virtual GPU (vGPU). This setup offers a balance of processing power, memory capacity, and advanced GPU acceleration—crucial factors for efficiently handling the intensive computations required by our QFNN-FFD framework.
@@ IV-B Accuracy and Loss Analysis
IV-C Local Client Validation Accuracy Analysis
IV-D Quantum Noise Analysis
@@ IV-E Comparison with Existing Works
Reference Precision Recall F1-Score Accuracy Privacy [17] 84 84.44 75.68 83.92 $\times$ [18] 96.1 79.5 86 94.5 $\times$ [19] 90 $\times$ QNN 93 94 94 93 $\times$ Our QFNN-FFD 95 96 95 95
@@
Our purpose is not to compare classical versus quantum methods but specifically to evaluate the performance of our framework against existing QML frameworks for fraud detection. Compared to the results in Table I, our QFNN-FFD outperforms other QML models applied to similar datasets, achieving superior performance metrics. These metrics include precision, recall, F1-score, and accuracy, where QFNN-FFD demonstrates comprehensive superiority across all fronts. This performance is a testament to the model’s efficacy and highlights its ability to effectively integrate complex quantum computations within an FL framework. Unlike the existing models [17, 18, 19], which focus solely on performance, QFNN-FFD additionally integrates a privacy-preserving FL approach. This ensures high detection accuracy of 95% and enhanced data privacy, establishing QFNN-FFD as a leading solution for secure and efficient fraud detection in fintech.
IV-F Ablation Study
To further analyze the effectiveness of the proposed QFNN-FFD, we conduct an ablation study comparing its performance against a standalone QNN to evaluate the impact of federated learning on accuracy, privacy preservation, and stability. Both models utilize identical architectures, and hyperparameters, with QNN trained centrally while QFNN-FFD operates across 15 federated clients. The results indicate that QFNN-FFD achieves an accuracy of 95% compared to 93% for QNN (see Table I), demonstrating that federated training does not compromise performance while enhancing privacy. Furthermore, QFNN-FFD stabilizes learning across clients, as seen in client-wise validation accuracy trends, mitigating inconsistencies observed in the standalone QNN. Unlike QNN, which requires centralized data aggregation, QFNN-FFD preserves privacy by ensuring data remains local to clients, achieving comparable or superior performance while addressing real-world privacy concerns. These findings validate that QFNN-FFD maintains strong fraud detection capabilities while integrating privacy-preserving mechanisms, making it a promising approach for secure financial applications.
IV-G Discussion